OpenStudy (anonymous):

A group of people spent exactly $100 on tickets to a show. Adult tickets cost $11 and child tickets cost $8. How many children were in the group?

OpenStudy (anonymous):

A*8 + B*11 = 100

OpenStudy (anonymous):

11x4=44 7x8=56 56+44=100

OpenStudy (anonymous):

7 kids 4 adults

OpenStudy (anonymous):

yeap 7kids n 4 adults

OpenStudy (anonymous):

Romero wrote an equation A*8+B*11 = 100. If were not given how many tickets are sold then we can do a guess and check method. If you plug in values for B you will find that when B = 4, then B*11 = $44. So if we take the original $100 and subtract $44 we get 56, which is divisible by 8.

OpenStudy (anonymous):

thanx you guys

OpenStudy (anonymous):

Where A and B have to be positive integers. The only solution for which satisfies the equation is 7 kids and 4 adults

OpenStudy (anonymous):

thanx